OpenStudy (anonymous):

Compare and contrast, in complete sentences, the difference between the products of x3x4 and (x3)4. Provide the expanded form and product of each.

OpenStudy (anonymous):

\[x^3x^4=x^7\]\[(x^3)^4=x^{12}\]

OpenStudy (anonymous):

x^3x^4 = (x*x*x)(x*x*x*x) = x^7 (x^3)^4 = (x^3)(x^3)(x^3)(x^3) = (x*x*x)(x*x*x)(x*x*x)(x*x*x) = x^12

OpenStudy (anonymous):

so x^3x^4 = x^3+4 = x^7 and (x^3)^4 = x^3*4 = x^12